# Altermagnetic Spin Precession and Spin Transistor

### v1

@@ Abstract
Altermagnets hold great potential for spintronic applications, yet their intrinsic spin dynamics and associated transport properties remain largely unexplored. Here, we investigate spin-resolved quantum transport in a multi-terminal setup based on a $d$-wave altermagnet. It is found that the altermagnetic spin splitting in momentum space induces an interesting spin precession in real space, giving rise to characteristic spin patterns. This altermagnetic spin precession manifests as a spatial modulation of the Hall voltage, whose oscillation period provides a direct measure of the spin-splitting strength. When the altermagnetism is electrically tunable, the proposed setup functions as a prototype for a highly efficient spin transistor. The key physical effects are shown to be robust against dephasing and crystalline warping. Our work not only identifies a fingerprint signature of altermagnets, offering a direct probe of the altermagnetic spin splitting, but also represents an important step toward bridging their fundamental physics with practical spintronic applications.
@@
Figure 1: (a) Schematic of the proposed setup, consisting of a central altermagnetic region connected to two longitudinal leads (1 and 2) and multiple transverse voltage probes. Top and bottom gates ($V_{tg}$ and $V_{bg}$) control the altermagnetic splitting, which is essential for the spin transistor. The $d$-wave altermagnetic Fermi surfaces are illustrated in the inset. (b) Altermagnetic spin precession in real space, with arrows indicating the local spin orientations. (c,d) Spatial distribution of spin component $S_{y}$ for nonequilibrium propagating electrons under (c) point injection and (d) line injection, with the injected spins polarized along the $x$ direction. The solid and dashed arrows highlight the corresponding spin modulation along the $x$ direction in (c) and (d). A square-root scale is applied to the colorbar to enhance the contrast around zero.
@@
Significant progress has been made in the search for altermagnetic materials, both theoretically and experimentally bai2024altermagnetism; song2025altermagnets. In particular, first-principles calculations have predicted hundreds of potential altermagnetic candidates bai2024altermagnetism; sodequist2024two; vsmejkal2022beyond; vsmejkal2022emerging; jiang2025metallic; zhang2025crystal; hu2025catalog; chen2024enumeration, and altermagnetic spin splitting has been experimentally observed through angle-resolved photoemission spectroscopy zhu2024observation; fedchenko2024observation; krempasky2024altermagnetic; ding2024large; lin2024observation; reimers2024direct; jiang2025metallic; zhang2025crystal; yang2025three; jungwirth2025symmetry. A more refined characterization of the altermagnetic Fermi surface, through quantum oscillation measurements li2025diagnosing and quasiparticle interference imaging hu2025quasiparticle; lou24prb; linder24prb, is warranted. The distinctive band structures of altermagnets can lead to a range of notable physical phenomena, including the anomalous Hall effect vsmejkal2020crystal; feng2022anomalous; vsmejkal2022anomalous; reichlova2024observation; gonzalez2023spontaneous; zhou2025manipulation; galindez2025revealing; leiviska2024anisotropy; li2025exploration; yu2025neel, giant magnetoresistance vsmejkal2022beyond; vsmejkal2022giant; li2025marginal, magnetic nonlinear Hall effect han2025discovery, spin-charge conversion karube2022observation; bai2023efficient; gonzalez2021efficient; jechumtal2025spin, interesting Josephson effects li2025spin; zhang2024finite, novel magnetoelectric coupling chen2025electrical; duan2025antiferroelectric; gu2025ferroelectric and Hall drag effects lin2025coulomb. While these advances highlight the great potential of altermagnets for spintronic applications, the electron spin dynamics and the resulting quantum transport properties remain poorly understood.
@@
In this work, we address this gap by investigating spin-resolved quantum transport in a multi-terminal device based on a two-dimensional $d$-wave altermagnet, as illustrated in Fig. 1(a). Owing to the altermagnetic spin splitting, electrons injected with spin polarization perpendicular to the Néel vector of the altermagnet undergo a characteristic spin precession [Fig. 1(b)], resulting in distinctive real-space spin patterns [Figs. 1(c,d)]. These spin patterns faithfully reflect the underlying altermagnetic spin splitting in momentum space and can be probed through the spatial oscillations of the Hall voltage. The oscillation period, in particular, is an analytical function of the altermagnetic spin splitting, providing a direct measure of this key physical parameter. The physical effects are robust against both dephasing and crystalline warping. Moreover, for electrically tunable altermagnets, the proposed setup naturally functions as a prototype for a high-efficiency spin transistor. Our results establish a fingerprint signature of altermagnets and pave the way for their device applications.
@@
Similar to the spin-orbit coupling effect, the altermagnetic spin splitting can be viewed as a momentum-resolved Zeeman field, leading to distinctive spin precession and, consequently, characteristic spin patterns in real space. To illustrate this, we first consider the local injection of electrons at the origin without loss of generality. The propagator from the origin to $\bm{r}$ takes the form SM
@@
The feature of altermagnetic spin precession becomes evident in the far-field limit $\bm{k}_{\sigma}\cdot\bm{r}\rightarrow\infty$, where the modified Bessel function exhibits the asymptotic behavior $K_{0}(-i\bm{k}_{\sigma}\cdot\bm{r})\rightarrow\sqrt{\frac{\pi}{2\bm{k}_{\sigma}\cdot\bm{r}}}e^{i(\bm{k}_{\sigma}\cdot\bm{r}+\frac{\pi}{4})}$ arfken2011mathematical. In this limit, the spatial spin distribution can be evaluated by the wave function as $(S_{x},S_{y})=(\psi^{\dagger}\sigma_{x}\psi,\psi^{\dagger}\sigma_{y}\psi)=A(\bm{r})\left[\cos(\Delta\bm{k}\cdot\bm{r}),~-\sin(\Delta\bm{k}\cdot\bm{r})\right]$, with $\Delta\bm{k}=(\Delta k_{x},\Delta k_{y})=\bm{k}_{\uparrow}-\bm{k}_{\downarrow}$ and $A=\left[(B^{2}-\alpha_{A}^{2})\sqrt{(\bm{k}_{\uparrow}\cdot\bm{r})(\bm{k}_{\downarrow}\cdot\bm{r})}/(8\pi^{3})\right]^{-1}$ the amplitude. Owing to the $d_{xy}$-symmetry of the altermagnetic spin splitting, electrons propagating along the $\pm\theta$ directions experience effective Zeeman fields of equal magnitude but opposite sign, resulting in opposite spin precessions shown in Fig. 1(b). As a result, the spatial distribution satisfies $S_{y}(x,y)=-S_{y}(x,-y)$, which can be verified in Fig. 1(c).
@@
Figure 2: (a) Transmission probabilities from lead 1 to the upper (blue dashed line) and lower (blue solid line) probes as a function of $x$. The red solid line represents the Hall voltage. (b) Hall voltage as a function of $x$ for different altermagnetic splitting strengths $\alpha_{A}$, with its zeros marked by green points. The gray dashed lines correspond to $\tilde{x}_{n}(\alpha_{A})$. (c,d) Same plots including the dephasing effect with $\Gamma_{v}=0.03$. The relevant parameters are $B=2$, $\mu=1$, $\alpha_{A}=0.8$, and $a=0.5$. The altermagnetic region has a size of $50\times 50$, and the hopping between the leads and the scattering region is $0.1t$.
Nonlocal spin-resolved transport.—We consider spin-resolved transport measurements in the Hall bar device as illustrated in Fig. 1(a). The setup consists of a square-shaped altermagnet connected to two longitudinal leads (leads 1 and 2), along with multiple transverse narrow leads designed to probe the spatially-resolved spin information. The transport properties are calculated based on the lattice model. Specifically, the $d_{xy}$-type altermagnet can be mapped onto a 2D square lattice as hallberg2025visualization
@@
The transmission probability from lead $q$ to $p$ is calculated by $T_{pq}=\mathrm{Tr}\left[\Gamma_{p}G^{R}\Gamma_{q}G^{A}\right]$, where the retarded and advanced Green’s functions are defined by $G^{R}=\left[G^{A}\right]^{\dagger}=\left[\omega-H_{\text{latt}}-\sum_{p}\Sigma_{p}^{R}\right]^{-1}$ and the linewidth function is given by $\Gamma_{p}=i\left(\Sigma^{R}_{p}-\Sigma^{A}_{p}\right)$. Both expressions contain the self-energies $\Sigma^{R}_{p}=\left[\Sigma^{A}_{p}\right]^{\dagger}=t_{p}^{2}G^{R}_{s,p}$ due to the coupling between lead $p$ and scattering region with strength $t_{p}$, where $G^{R}_{s,p}$ is the surface Green’s function of lead $p$ sancho1985highly. In our calculations, we consider current flowing between leads 1 and 2, with $I_{1}=-I_{2}$. The upper ($u$) and lower ($l$) leads, labeled by their positions $x$, serve as voltage probes that measure the transverse voltages $V_{u,l}(x)$ under the condition of vanishing current. The voltages at all leads, including $V_{1,2}$ and $V_{u,l}(x)$, are then solved by Eq. (7).
@@
In Fig. 2(a), we plot the spin-resolved transmission probabilities $T_{u}(x)$ and $T_{l}(x)$ from lead 1 to the upper and lower voltage probes as a function of their position $x$. The altermagnetic spin precession induces a periodic modulation of the spin overlap between the propagating electrons and the voltage probes, resulting in spatial oscillations of the transmission probabilities. By comparing with the spin textures near $y=\pm W/2$ in Fig. 1(d), one can see that the modulation of transmission probabilities faithfully reveals the underlying spin precession. In particular, electrons reaching the upper and lower probes experience opposite spin precessions, such that their spin overlaps exhibit opposite spatial modulations. This explains the nearly $\pi$ phase shift between the oscillations of $T_{u}$ and $T_{l}$ in Figs. 2(a,c). These spatial modulations of the transmission probabilities are inherited by the Hall voltage, defined as $V_{H}(x)=V_{u}(x)-V_{l}(x)$; see Fig. 2(a). The zeros of the Hall voltage occur at positions where $T_{u}=T_{l}$, corresponding to points where the spin precession angle equals $n\pi$ for both upward and downward propagations, such that their spin overlaps with the corresponding probes are identical. Interestingly, these zero points are closely related to the solutions in Eq. (4). This correspondence is illustrated in Fig. 2(b), where the Hall voltage modulations for different altermagnetic splittings are plotted, and the zero points (green dots) are compared with $\tilde{x}_{n}=1.3x_{n}$ (dashed lines), showing quantitative agreement. Here, the factor of $1.3$ arises from the reflection at the upper and lower boundaries. Since $\mathcal{T}_{x}$ is determined by the strength $\alpha_{A}$ of altermagnetic spin splitting, our scheme provides a straightforward approach for its measurement.
@@
Dephasing and warping effects.—The above calculations assume full phase coherence. In reality, however, dephasing and spin relaxation may occur due to inelastic scattering and interactions with magnetic impurities wu2010spin; jiang2009topological. Due to the altermagnetic coupling between spin and momentum, these two effects generally occur simultaneously. To incorporate these effects, we employ Büttiker’s approach, modeling dephasing by attaching virtual voltage probes buttiker1986role; golizadeh2007nonequilibrium; zhou2022transport. Each lattice site i in the altermagnetic region is coupled to a probe represented by the self-energy $\Sigma_{\textbf{i}}^{R}=-i\Gamma_{v}/2$, where $\Gamma_{v}$ characterizes the dephasing strength, i.e., the probability for electrons to escape into the virtual probes and lose their phase information. Mathematically, the self-energies of both real and virtual leads enter the Green’s function on an equal footing, i.e., $G^{R}=\left[\omega-H_{\text{latt}}-\sum_{p}\Sigma_{p}^{R}-\sum_{\textbf{i}}\Sigma_{\textbf{i}}^{R}\right]^{-1}$, from which the transmission matrix taking into account both real and virtual leads can be obtained. Using the same current configuration as in the case without dephasing, and imposing the condition of zero current through all voltage probes, the Hall voltages are solved by Eq. (7). In our calculations, we set $\Gamma_{v}=0.03$, which corresponds to a coherent-to-total current ratio of 41%.
@@
Figure 3: Transmission probabilities and Hall voltage as a function of $\alpha_{A}$ at $x=12.5$. All other parameters are the same as in Fig. 2.
Altermagnetic spin transistor.—Altermagnets have shown great potential for spintronic applications owing to their strong spin splitting, which is an order of magnitude larger than that induced by spin–orbit coupling vsmejkal2022emerging, thereby enabling spintronic devices with higher efficiency and smaller sizes. When altermagnetic spin splitting can be controlled electrically wang2024electric, the proposed device can function as a spin transistor datta1990electronic. We have seen that, for a fixed altermagnetic spin splitting $\alpha_{A}$, the spin precession gives rise to a location-dependent Hall voltage $V_{H}$, with a period determined by $\alpha_{A}$ through Eq. (4). Consequently, tuning $\alpha_{A}$ is expected to modulate both the spin precession and therefore $V_{H}$, thereby realizing a spin transistor. In Fig. 3, we show the transmissions $T_{u,l}$ and $V_{H}$ as functions of $\alpha_{A}$, where only a single pair of Hall probes is involved. The results show that the voltage drop can be efficiently tuned, enabling clear switching with a high on/off ratio.
@@
Discussions.— It is found that the altermagnetic spin splitting induces a distinctive spin precession whose sense and rate depend strongly on the electron propagation direction, thereby establishing a clear correspondence between the spin structures in real and momentum spaces. This precession leads to a spatial modulation of the Hall voltage, whose period directly reflects the magnitude of the altermagnetic spin splitting, allowing its precise determination in the transport setup. The predicted effect and its signature are shown to be robust against both dephasing and crystalline warping. For a given material, the spatial modulation of the spin texture depends on several parameters such as the altermagnetic spin splitting, effective mass, and chemical potential. To resolve the signature with high fidelity, the width of the Hall voltage probes should be smaller than, or at most comparable to, the modulation period. Although we focus here on the $d$-wave altermagnet, similar analyses can be carried out for other symmetries vsmejkal2022emerging; vsmejkal2022beyond, where intriguing spin precession is likewise expected. As this spin precession represents an intrinsic physical property of altermagnets, it has broad implications for their spintronic applications, such as generating complex spin textures, controlling magnetization dynamics, and manipulating domain walls and skyrmions manchon2019current; obata2008current.

### v2

@@ Abstract
Altermagnets hold great potential for spintronic applications, yet their intrinsic spin dynamics and associated transport properties remain largely unexplored. Here, we investigate spin-resolved quantum transport in a multi-terminal setup based on a $d$-wave altermagnet. It is found that the altermagnetic spin splitting in momentum space induces an interesting spin precession in two-dimensional real space, giving rise to characteristic spin patterns. This altermagnetic spin precession manifests as a spatial modulation of the transverse Hall-like voltage, whose oscillation period provides a direct measure of the spin-splitting strength. When the altermagnetism is electrically tunable, the proposed setup functions as a prototype for a highly efficient spin transistor. The key physical effects are shown to be robust against dephasing and crystalline warping. Our work not only identifies a fingerprint signature of altermagnets, offering a direct probe of the altermagnetic spin splitting, but also represents an important step toward bridging their fundamental physics with practical spintronic applications.
@@
Figure 1: (a) Schematic of the proposed setup, consisting of a central altermagnetic region connected to two longitudinal leads (1 and 2) and multiple transverse voltage probes. Top and bottom gates ($V_{tg}$ and $V_{bg}$) control the altermagnetic splitting, which is essential for the spin transistor. The $d$-wave altermagnetic Fermi surfaces are illustrated in the inset. (b) Altermagnetic spin precession in real space, with arrows indicating the local spin orientations. (c,d) Spatial distribution of spin component $S_{y}$ for nonequilibrium propagating electrons under (c) point injection and (d) line injection, with the injected spins polarized along the $x$ direction. The solid and dashed arrows highlight the corresponding spin modulation along the $x$ direction in (c) and (d). A square-root scale is applied to the colorbar to enhance the contrast around zero. Relevant parameters: $B=0.2\mathrm{~eV}\cdot\mathrm{nm}^{2},~\alpha_{A}=0.08\mathrm{~eV}\cdot\mathrm{nm}^{2},~\mu=0.1\mathrm{~eV}$.
@@
Significant progress has been made in the search for altermagnetic materials, both theoretically and experimentally bai2024altermagnetism ; song2025altermagnets . In particular, first-principles calculations have predicted hundreds of potential altermagnetic candidates bai2024altermagnetism ; sodequist2024two ; vsmejkal2022beyond ; vsmejkal2022emerging ; jiang2025metallic ; zhang2025crystal ; hu2025catalog ; chen2024enumeration , and altermagnetic spin splitting has been experimentally observed through angle-resolved photoemission spectroscopy zhu2024observation ; fedchenko2024observation ; krempasky2024altermagnetic ; ding2024large ; lin2024observation ; reimers2024direct ; jiang2025metallic ; zhang2025crystal ; yang2025three ; jungwirth2025symmetry . A more refined characterization of the altermagnetic Fermi surface, through quantum oscillation measurements li2025diagnosing and quasiparticle interference imaging hu2025quasiparticle ; lou24prb ; linder24prb , is warranted. The distinctive band structures of altermagnets can lead to a range of notable physical phenomena, including the anomalous Hall effect vsmejkal2020crystal ; feng2022anomalous ; vsmejkal2022anomalous ; reichlova2024observation ; gonzalez2023spontaneous ; zhou2025manipulation ; galindez2025revealing ; leiviska2024anisotropy ; li2025exploration ; yu2025neel , giant magnetoresistance vsmejkal2022beyond ; vsmejkal2022giant ; li2025marginal , magnetic nonlinear Hall effect han2025discovery , spin-charge conversion karube2022observation ; bai2023efficient ; gonzalez2021efficient ; jechumtal2025spin , interesting Josephson effects li2025spin ; zhang2024finite , novel magnetoelectric coupling chen2025electrical ; duan2025antiferroelectric ; gu2025ferroelectric and Hall drag effects lin2025coulomb , and other interesting transport properties vakili2025spin ; kokkeler2025quantum ; zhang2025electrical ; zarzuela2025transport . While these advances highlight the great potential of altermagnets for spintronic applications jungwirth2025altermagnetic , their characteristic spin dynamics and the resulting quantum transport properties in multi-terminal measurements remain poorly understood.
@@
In this work, we address this gap by investigating spin-resolved quantum transport in a multi-terminal device based on a two-dimensional $d$-wave altermagnet, as illustrated in Fig. 1(a). Owing to the altermagnetic spin splitting, electrons injected with spin polarization perpendicular to the Néel vector of the altermagnet undergo a characteristic spin precession [Fig. 1(b)], resulting in distinctive real-space spin patterns [Figs. 1(c,d)]. These spin patterns faithfully reflect the underlying altermagnetic spin splitting in momentum space and can be probed through the spatial oscillations of the transverse Hall-like voltage. The oscillation period, in particular, is an analytical function of the altermagnetic spin splitting, providing a direct measure of this key physical parameter. The physical effects are robust against both dephasing and crystalline warping. Moreover, for electrically tunable altermagnets, the proposed setup naturally functions as a prototype for a high-efficiency spin transistor. Our results establish a fingerprint signature of altermagnets and pave the way for their device applications.
@@
The altermagnetic spin splitting can be viewed as a momentum-resolved Zeeman field, leading to distinctive spin precession and, consequently, characteristic spin patterns in real space. To illustrate this, we first consider the local injection of electrons at the origin without loss of generality. The propagator from the origin to $\bm{r}$ takes the form SM
@@
The feature of altermagnetic spin precession becomes evident in the far-field limit $\bm{k}_{\sigma}\cdot\bm{r}\rightarrow\infty$, where the modified Bessel function exhibits the asymptotic behavior $K_{0}(-i\bm{k}_{\sigma}\cdot\bm{r})\rightarrow\sqrt{\frac{\pi}{2\bm{k}_{\sigma}\cdot\bm{r}}}e^{i(\bm{k}_{\sigma}\cdot\bm{r}+\frac{\pi}{4})}$ arfken2011mathematical . In this limit, the spatial spin distribution can be evaluated by the wave function as $(S_{x},S_{y})=(\psi^{\dagger}\sigma_{x}\psi,\psi^{\dagger}\sigma_{y}\psi)=A(\bm{r})\left[\cos(\Delta\bm{k}\cdot\bm{r}),~-\sin(\Delta\bm{k}\cdot\bm{r})\right]$, with $\Delta\bm{k}=(\Delta k_{x},\Delta k_{y})=\bm{k}_{\uparrow}-\bm{k}_{\downarrow}$ and $A=\left[(B^{2}-\alpha_{A}^{2})\sqrt{(\bm{k}_{\uparrow}\cdot\bm{r})(\bm{k}_{\downarrow}\cdot\bm{r})}/(8\pi^{3})\right]^{-1}$ the amplitude. The distinctive spin precession is visualized in Fig. 1(b). Specifically, as $\theta$ increases from 0 to $\pi/4$, the spin precession rate gradually increases from zero (vanishing precession) to its maximum; and with further increasing $\theta$, the spin precession rate begins to decrease monotonically, until it reaches zero at $\theta=\pi/2$. Furthermore, the spin precession is mirror symmetric relative to the $x$ axis, i.e., the precessions for $\theta$ and $\theta$ have the same rate but opposite directions. As a result, the spatial spin distribution satisfies $S_{y}(x,y)=-S_{y}(x,-y)$, which can be verified in Fig. 1(c).
@@
Figure 2: (a) Transmission probabilities from lead 1 to the upper (blue dashed line) and lower (blue solid line) probes and Hall voltage (red solid line) as a function of $x$ with $\alpha_{A}=0.08\mathrm{~eV}\cdot\mathrm{nm}^{2}$. (b) A waterfall-like plot of Hall voltage as a function of $x$ for different altermagnetic splitting $\alpha_{A}$, with its zeros marked by green points. The gray dashed lines correspond to $\tilde{x}_{n}(\alpha_{A})$. (c, d) Same plots including the dephasing effect with $\Gamma_{v}=3\mathrm{~meV}$. The size of the scattering region is $50\mathrm{~nm}\times 50\mathrm{~nm}$, and the lattice constant is $a=0.5\mathrm{~nm}$. The interface hopping is set to $t$ for leads 1 and 2, and to $0.1t$ for the transverse leads. All other parameters are the same as those in Fig. 1, which corresponds to the spatial period $\tilde{\mathcal{T}}_{x}=4.67\text{ nm}$.
Nonlocal spin-resolved transport.—We consider spin-resolved transport measurements in the Hall bar device as illustrated in Fig. 1(a). The setup consists of a square-shaped altermagnet connected to two longitudinal leads (leads 1 and 2), along with multiple transverse narrow leads designed to probe the spatially-resolved spin information. The transport properties are calculated based on the lattice model. Specifically, the $d_{xy}$-type altermagnet can be mapped onto a 2D square lattice with lattice constant $a$ as hallberg2025visualization
@@
[…]
@@
In Fig. 2(a), we plot the transmission probabilities $T_{u}(x)$ and $T_{l}(x)$ from lead 1 to the upper and lower voltage probes as a function of their position $x$. The altermagnetic spin precession induces a periodic modulation of the spin overlap between the propagating electrons and the voltage probes, resulting in spatial oscillations of the transmission probabilities. By comparing with the spin textures near $y=\pm W/2$ in Fig. 1(d), one can see that the modulation of transmission probabilities faithfully reveals the underlying spin precession. In particular, electrons reaching the upper and lower probes experience opposite spin precessions, such that their spin overlaps exhibit opposite spatial modulations. This explains the nearly $\pi$ phase shift between the oscillations of $T_{u}$ and $T_{l}$ in Figs. 2(a,c).
These spatial modulations of the transmission probabilities are inherited by the Hall voltage, defined as $V_{H}(x)=V_{u}(x)-V_{l}(x)$; see Fig. 2(a). Specifically, the potential at the voltage probe $p$ is given by $V_{p}=\frac{\sum_{q\neq p}T_{pq}V_{q}}{\sum_{q\neq p}T_{pq}}$ datta1997electronic . In the present setup, this yields $V_{u(l)}(x)\simeq\frac{T_{u(l)}V_{1}}{T_{u(l)}+T_{u(l)2}}$, where $T_{u(l)2}$ denotes the transmission from lead 2 to the upper (lower) voltage probe. We have set $V_{2}=0$ without loss of generality and neglected the small transmissions between voltage probes at different positions $x$. The Hall voltage vanishes at positions $\tilde{x}_{n}$ determined by $\tilde{S}_{y}(\tilde{x}_{n},\pm W/2)=0$. At these points, the local spin overlaps with the upper and lower probes are identical, leading to $T_{u}=T_{l}$. Moreover, the symmetry of the system ensures $T_{u2}=T_{l2}$ at $\tilde{x}_{n}$, which directly yields $V_{H}(\tilde{x}_{n})=0$. This correspondence is illustrated in Fig. 2(b), where the Hall voltage modulations for different altermagnetic splittings are plotted, and the zero points (green dots) are compared with $\tilde{x}_{n}=n\tilde{\mathcal{T}}_{x}=1.3x_{n}$ (dashed lines), showing quantitative agreement. Here, the factor of $1.3$ arises from reflections at the upper and lower boundaries present in the transport calculations but absent in Fig. 1(d). Since $\mathcal{T}_{x}$ is determined by the strength $\alpha_{A}$ of altermagnetic spin splitting, our scheme provides a direct and experimentally accessible means for its measurement. The altermagnetic spin splitting spans a wide range, from $10~\mathrm{meV}$ to $1400~\mathrm{meV}$ vsmejkal2022emerging , implying that the corresponding spin-precession period can vary from a few nanometers to several tens of nanometers. For a representative current of $I_{1}=1~\mathrm{\mu A}$, the corresponding Hall-voltage range in Figs. 2(a,c) is $0.13,0.13]~\mathrm{mV}$, which is well within experimental reach. The above analysis indicates that the predicted spin precession can be observed in intrinsic altermagnets such as KRu4O8 and V2OS.
@@
Dephasing and warping effects.—The above calculations assume full phase coherence. In reality, however, dephasing and spin relaxation may occur due to inelastic scattering and interactions with magnetic impurities wu2010spin ; jiang2009topological . Due to the altermagnetic coupling between spin and momentum, these two effects generally occur simultaneously. To incorporate these effects, we employ Büttiker’s approach, modeling dephasing by attaching virtual voltage probes buttiker1986role ; golizadeh2007nonequilibrium ; zhou2022transport . Each lattice site i in the altermagnetic region is coupled to a probe represented by the self-energy $\Sigma_{\textbf{i}}^{R}=-i\Gamma_{v}/2$, where $\Gamma_{v}$ characterizes the dephasing strength, i.e., the probability for electrons to escape into the virtual probes and lose their phase information. Mathematically, the self-energies of both real and virtual leads enter the Green’s function on an equal footing, i.e., $G^{R}=\left[\omega-H_{\text{latt}}-\sum_{p}\Sigma_{p}^{R}-\sum_{\textbf{i}}\Sigma_{\textbf{i}}^{R}\right]^{-1}$, from which the transmission matrix taking into account both real and virtual leads can be obtained. Using the same current configuration as in the case without dephasing, and imposing the condition of zero current through all voltage probes, the Hall voltages are solved by Eq. (7). In our calculations, we set $\Gamma_{v}=3~\mathrm{meV}$, which corresponds to a coherent-to-total current ratio of 41%.
@@
Figure 3: Transmission probabilities and Hall voltage as a function of $\alpha_{A}$ at $x=12.5\mathrm{~nm}$. All other parameters are the same as in Fig. 1.
Altermagnetic spin transistor.—Altermagnets have shown great potential for spintronic applications owing to their strong spin splitting, which is an order of magnitude larger than that induced by spin-orbit coupling vsmejkal2022emerging , thereby enabling spintronic devices with higher efficiency and smaller sizes. When altermagnetic spin splitting can be controlled electrically wang2024electric , the proposed device can function as a spin transistor datta1990electronic . We have seen that, for a fixed altermagnetic spin splitting $\alpha_{A}$, the spin precession gives rise to a location-dependent Hall voltage $V_{H}$, with a period determined by $\alpha_{A}$ through Eq. (4). Consequently, tuning $\alpha_{A}$ is expected to modulate both the spin precession and therefore $V_{H}$, thereby realizing a spin transistor. In Fig. 3, we show the transmissions $T_{u,l}$ and $V_{H}$ as functions of $\alpha_{A}$, where only a single pair of Hall probes is involved. The results show that the voltage drop can be efficiently tuned, enabling clear switching with a high on/off ratio.
@@
It is instructive to highlight the differences between the altermagnetic spin transistor and the Datta-Das spin transistor datta1990electronic , which relies on gate-controlled spin-orbit coupling. Although both scenarios share a collinear spin structure, their distinct dimensionalities and momentum dependences give rise to qualitatively different spin dynamics and physical outcomes. In particular, Datta-Das spin precession manifests itself as oscillations in a longitudinal transport signal, whereas altermagnetic spin precession is revealed through oscillations in a transverse transport response. The larger spin splitting intrinsic to altermagnets enables high-frequency spin manipulation within a more compact device footprint, therefore enhancing robustness against dephasing. Furthermore, the broader range of candidate altermagnetic materials offers a promising route to overcome the long-standing spin-injection limitation due to the conductivity mismatch at interfaces vzutic2004spintronics .
Discussions.— The altermagnetic spin precession and transport properties are highly anisotropic and depend sensitively on the relative orientation between the applied bias and the crystallographic direction. In contrast to transport along the nodal direction, i.e., the $x$ direction for the $d_{xy}$ Fermi-surface configuration, transport along the anti-nodal direction yields a vanishing $V_{H}$; see the Appendix for details. This pronounced anisotropy further highlights the distinctive characteristics of altermagnetism. Although we focus here on the $d$-wave altermagnet, similar analyses can be carried out for other symmetries vsmejkal2022emerging ; vsmejkal2022beyond , where intriguing spin precession is likewise expected. As this spin precession represents an intrinsic physical property of altermagnets, it has broad implications for their spintronic applications, such as generating complex spin textures, controlling magnetization dynamics, and manipulating domain walls and skyrmions manchon2019current ; obata2008current .